1. During the access today I checked the two BB shower channels that were reading zero current. The cables for these channels were not properly connected(this could have happened during the HV crate swap yesterday) so I unplugged and plugged back them back in and it works fine now.2. For the scalers on LHRS. I sent another copy of the T2 and T7 to LHRS on a different cable and this copy goes to scaler right now. We'll see if this works when the beam is back.